OpenStudy (anonymous):

Find the solution to the following set of equations: 3x + 10y = 14 4x + 5y =2 I need to solve this problem using elimination. Can you please help me solve this problem?

OpenStudy (baru):

multiply second equation by '2' on both sides. what do you get?

OpenStudy (anonymous):

8x + 10y = 4

OpenStudy (baru):

yes exactly :)

OpenStudy (anonymous):

Ok, but i though the x's needed to have the same number, like 12.

OpenStudy (baru):

now subtract equation 1 from equation 2

OpenStudy (anonymous):

o 10 y ok i see

OpenStudy (baru):

yes exactly :) i'm doing it to y instead of x

OpenStudy (anonymous):

o ok. I can get that one then. What about this one. 2x + 3y = 0 and 5x - 4y = 23?

OpenStudy (anonymous):

Would I do this one the way i tried to do the other one?

OpenStudy (baru):

yes, you need to recognize with what number you need to multiply the two equations so that the 'x' or the y's match

OpenStudy (baru):

in this case, multiply equation 1 by 5 and equation 2 by 2

OpenStudy (anonymous):

ok so i'd multiply the first equation by 5 and the second equation by 2? Right?

OpenStudy (anonymous):

o ok haha sorry

OpenStudy (anonymous):

then i'd subract the equations from each other correct?

OpenStudy (anonymous):

Because the x's are both positive.

OpenStudy (baru):

yes, you are correct :)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

So I got y=2 and x=-3

OpenStudy (baru):

yep thats right

OpenStudy (anonymous):

Ok thank you. This makes much more sense to me.

OpenStudy (baru):

sure, no prob :)